3x + 2y =8 and y= X-1, then x=

Found 2 solutions by solver91311, zkiya_:
Answer by solver91311(24713)



No solution. That is because and are NOT the same thing.

Now, if what you REALLY meant was that your system of linear equations is:





That is a horse of a completely different color. Since the second equation gives you an expression in that is equal to, and that means "the same as", , you can replace anywhere you see it with that expression in , namely , thus:



Now all you have to do is solve the single variable equation for

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism



Answer by zkiya_(6)

3x + 2y = 8
y = x-1
3x + 2(x-1) = 8
3x + 2x - 2 = 8
5x = 10
x = 2